# Covering Multiple Objectives with a Small Set of Solutions Using Bayesian Optimization

### v1

@@ 3.1.2 Greedy Approximation of $S^{*}_{D_{s}}$
Since constructing $S^{*}_{D_{s}}$ is NP-hard, we use an approximate construction of $S^{*}_{D_{s}}$ on each step of optimization $s$. We present Algorithm 1 with $O(nKT)$ running time that achieves a constant $(1-\frac{1}{e})$ approximation factor.
@@
0:  Dataset $D_{s}$ with observed values $f_{t}(\mathbf{x})$ for all $t\in\{1,\ldots,T\}$ and $\mathbf{x}\in D_{s}$, covering set size $K$.
1:  Initialize $A\leftarrow\emptyset$. {Start with an empty set.}
2:  Compute the initial coverage score for $A$: $c(A)\leftarrow 0$.
3:  for $k=1$ to $K$ do
4:     Initialize $\mathbf{x}_{\text{best}}\leftarrow\text{None}$ and $\Delta_{\text{best}}\leftarrow-\infty$.
5:     for $\mathbf{x}\in D_{s}\setminus A$ do
6:        Compute the marginal coverage of adding $\mathbf{x}$ to $A$:
@@
7:        if $\Delta c>\Delta_{\text{best}}$ then
8:           $\mathbf{x}_{\text{best}}\leftarrow\mathbf{x}$ and $\Delta_{\text{best}}\leftarrow\Delta c$.
9:        end if
10:     end for
11:     Update $A\leftarrow A\cup\{\mathbf{x}_{\text{best}}\}$, $c(A)\leftarrow c(A)+\Delta_{\text{best}}$.
12:  end for
13:  Output: $A^{*}_{D_{s}}\leftarrow A$.
Theorem 3.2 (Algorithm 1 provides a $(1-\frac{1}{e})$-Approximation of $S^{*}_{D_{s}}$).
Algorithm 1 provides a $(1-\frac{1}{e})$-Approximation of $S^{*}_{D_{s}}$. Formally, for the approximate covering set $A^{*}_{D_{s}}$ output by Algorithm 1:
$c(A^{*}_{D_{s}})\geq\left(1-\frac{1}{e}\right)c(S^{*}_{D_{s}})$
@@
The algorithm iterates $K$ times to construct the covering set. In each iteration, it evaluates at most $n$ candidate points, and for each candidate, it computes the incremental coverage score by iterating over $T$ objectives. The runtime is thus $O(K\cdot n\cdot T)$. For practical applications where we can assume relatively small $K$ and $T$, the runtime is approximately $O(n)$.
@@
This follows from the fact that we reduced from Max $k$-cover, for which a better approximation ratio is not practically achievable unless $P=NP$ (Feige, 1998). ∎
3.1.3 MOCOBO Trust Region Dynamics
After each optimization step $s$, we center each trust region $\mathcal{T}_{k}$ on the corresponding point $\mathbf{x}^{*(s)}_{k}$ in the best observed covering set $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\right\}$ found so far, as defined in Equation 4. As in the original TuRBO algorithm, each trust region $\mathcal{T}_{k}$ has success and failure counters that dictate the size of the trust region. For MOCOBO, we count a success for trust region $\mathcal{T}_{k}$ whenever $\mathcal{T}_{k}$ proposes a candidate on step $s$ that improves upon the best coverage score and is included in $S^{*}_{D_{s+1}}$.
@@ 3.1.4 Extending ECI to the Batch Acquisition Setting (q-ECI)
In the batch acquisition setting, we select a batch of $q>1$ candidates for evaluation from each trust region. Following recent work on batch expected improvement (Wilson et al., 2017; Wang et al., 2019), we define q-ECI, a natural extension of ECI to the batch setting:
$q\text{-ECI}(\mathbf{X})=\\ \mathbb{E}_{p(\mathbf{Y}|\mathbf{X},D)}\left[\max(0,\max_{r=1,\dots,q}c(S^{*}_{D_{s}\cup\left\{(\mathbf{x}_{r},\mathbf{y}_{r})\right\}})-c(S^{*}_{D_{s}}))\right]$ (7)
q-ECI gives the expected improvement in the coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. However, the resulting Monte Carlo expectation would require $O(q\times m)$ evaluations of Algorithm 1. We therefore adopt a more approximate batching strategy for practical use with large $q$ (see Section A.3 for a discussion of q-ECI intractability and our approximation).
@@ 4 Experiments
We implement MOCOBO using BoTorch (Balandat et al., 2020) and GPyTorch (Gardner et al., 2018). Code to reproduce all results in the paper will be made publicly available on github after the paper is accepted, with a link to the git repository provided here placeholder. We use an acquisition batch size of $20$ for all tasks and across all methods compared. Further implementation details are provided in subsection A.4. For our structured optimization problems, we apply MOCOBO in the latent space of a pre-trained VAE on which we perform regular end-to-end updates with the surrogate model during optimization (Maus et al., 2022).
@@
In Figure 3, we plot the best coverage score $c$ Equation 1 obtained by the $K$ best covering solutions found so far after a certain number of function evaluations on each task. Since baseline methods are not designed to optimize coverage directly, instead aiming to find a single solution or a set of solutions to optimize the $T$ objectives, we plot the best coverage score $c$ obtained by the $K$ best covering solutions found by the method. All plots show mean coverage scores averaged over $20$ replications of each method, and show standard errors.
4.1 Baselines
In all plots, we compare MOCOBO against TuRBO, ROBOT, and MORBO ((Eriksson et al., 2019; Maus et al., 2023; Daulton et al., 2021)). For each method, we compare to the performance of the best set of $K<T$ solutions found.
@@ 4.1 Baselines
In the case of the two structured optimization tasks (molecule and peptide design), we apply LOL-BO and ROBOT as described by design for these problem settings. For MORBO, we take the straightforward approach of applying MORBO directly in the continuous latent space of the VAE model without other adaptations. The same pre-trained VAE model is used across methods compared.
Extending baselines to the coverage BO setting.
As TuRBO, LOL-BO, and ROBOT target single-objective optimization, we conduct $T$ independent runs to optimize each of the $T$ objectives for the multi-objective task. We use all solutions gathered from the $T$ runs to compute the best covering set of $K$ solutions found by each method.
Unlike LOL-BO and TuRBO which seek a single best solution for a given objective, a single run of ROBOT seeks a set of $M$ solutions that are pairwise diverse. We run ROBOT with $M=K$ so that each independent run of ROBOT seeks $M=K$ diverse solutions. The aggregate result of the $T$ independent runs for ROBOT is thus $K$ diverse solutions for each of the $T$ objectives. We compare to the best covering $K$ solutions from among those $K*T$ solutions. See Section A.4.4 for more details on diversity constraints used by ROBOT and the associated diversity hyperparameters used to run ROBOT on each task.
MORBO is a multi-objective optimization method and can thus be applied directly to each multi-objective optimization task. For each run of MORBO, we compare to the best covering set of $K$ solutions found among all solutions proposed by the run.
@@ 4.2 Tasks
We evaluate MOCOBO on two variations of the peptide design task: “template free” (TF) and “template constrained” (TC). For TF, we allow the optimizer to propose any sequence of amino acids. For TC, we add a constraint that any sequence proposed by the optimizer must have a minimum of $75$ percent sequence similarity to at least one of the $10$ template amino acid sequences in Table A.6. These $10$ templates were mined from extinct organisms and selected by Wan et al. (2024). The motivation of the template constrained task is to design peptides specifically likely to evade antibiotic resistance by producing “extinct-like” peptides that bacteria have not encountered in nature in thousands of years. We handle the optimization constraint by adapting techniques from SCBO (Eriksson & Poloczek, 2021).
@@
Ranolazine is a drug used to treat chest pain. The original Ranolazine MPO task from the Guacamol benchmark suite of molecular design tasks (Brown et al., 2019) aims to design an alternative to this drug: a molecule with a high fingerprint similarity to Ranolaize that includes fluorine. We extend this task to the multi-objective optimization setting by searching for alternatives to Ranolazine that include $T=6$ reactive nonmetal elements not found in Ranolazine: fluorine, chlorine, bromine, selenium, sulfur, and phosphorus. We aim to cover the $T=6$ objectives with $K=3$ molecules. We use the SELFIES-VAE introduced by Maus et al. (2022) to map the molecular space to a continuous $256$ dimensional space.
@@
The rover trajectory optimization task introduced by Wang et al. (2018) consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a fixed set of obstacles. To frame this as a multi-objective optimization task, we design $T$ unique obstacle courses for the rover to navigate. The obstacle courses are designed such that no single policy can successfully navigate all courses. We seek $K<T$ policies so that at least one policy enables the rover to avoid obstacles in each course. We evaluate on three unique instances with varying numbers of obstacle courses. For the instances of this task with $T=4$ and $T=8$, we seek to cover the objectives with $K=2$ solutions. For the $T=12$ instance, we seek to cover the objectives with $K=4$ solutions.
@@
In high dynamic range (HDR) images, some pixels (often associated with light sources) can dominate overall contrast, requiring adjustments to reveal detail in low-contrast areas, a problem known as tone mapping (Reinhard et al., 2005, Section 6). Tone mapping algorithms involve various tunable parameters, resulting in a high-dimensional optimization problem of subjectively perceived quality. Because commonly used metrics are only correlated with subjective quality, tuning such parameters has been done by trial and error and human-in-the-loop type schemes (Lischinski et al., 2006), including preferential BO-based approaches (Koyama et al., 2017, 2020).
We seek a covering set of $K=4$ solutions to optimize a set of $T=7$ image aesthetic (IAA) and quality (IQA) assessment metrics from the pyiqa library (Chen & Mo, 2022) (see metrics listed in Table A.3). Our practical goal is that, while we do not know a priori which metric is best for a particular image, covering all metrics may result in at least one high quality image. We use the “Stanford Memorial Church” Debevec & Malik (1997) and “desk lamp” Čadík (2008) benchmark images. We optimize over the $13$-dimensional parameter space of an established tone mapping pipeline. See Section A.5.1 for details.
@@ 4.3 Optimization Results
In Figure 3, we provide optimization results comparing MOCOBO to the baselines discussed above on all tasks. The results show that MOCOBO finds sets of $K$ solutions that achieve higher coverage scores across tasks.
The “T Individually Optimized Solutions” baseline appears as a horizontal dotted line in all plots of Figure 3, representing the average coverage score of $T$ individually optimized solutions, serving as an approximation of the best possible performance without the constraint of a limited $K<T$ solution set. Results in Figure 3 demonstrate that the smaller set of $K<T$ solutions identified by MOCOBO nearly equals the performance of the complete set of $T$ individually optimized solutions. This result depends on using domain knowledge to choose $K$ large enough to achieve it. Results in Figure 3 show that matching the performance of $T$ optimized solutions is possible with some values of $K\ll T$.
Table 1: Best set of $K=4$ peptides found by one run of MOCOBO for the “template free” peptide design task. We provide the APEX model’s predicted MIC for each sequence on each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.5. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TF1” in Figure 4 and Figure A.1 for in vitro MICs for this set of $K=4$ peptides. See Table A.1 for an analogous result on the “template constrained” peptide design task.
Figure 4: In vitro results for the two best “template free” (TF1, TF2) and two best “template constrained” (TC1, TC2) runs of MOCOBO for the peptide design task. Columns are the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target pathogenic bacteria B1,,$\ldots$,, … , B11 listed in Table A.5. (-) and (+) indicate Gram negative and Gram positive respectively. TF1 and TC1 correspond to the single runs of MOCOBO shown in Table 1 and Table A.1 respectively. Methods used to obtain in vitro MICs are provided in Section A.1.
Figure 5: (Four Leftmost Panels) $T=4$ varied obstacle courses as multiple objectives for the rover task and $K=2$ covering trajectories (blue, magenta) found by MOCOBO. The first trajectory (magenta) successfully navigates obstacle courses 2 and 3 (second and third panels from the left). The second trajectory (blue) navigates obstacle courses 1 and 4 (first and fourth panels from the left). (Middle Panel) Initial hdr desk image for the desk variation of the “image tone mapping” task. (Four Rightmost Panels) Images obtained by transforming the hdr desk image using the best covering set of $K=4$ solutions found by a single run of MOCOBO.
@@
In Table 1, we provide the $K=4$ peptides found by one run of MOCOBO for the “template free” peptide design task. For each of the four peptides, we provide the APEX model’s predicted MIC for each of the $11$ target bacteria. MIC values $\leq 16$ $\mu$ mol $L^{-1}$ are considered to be “highly active” against the target pathogenic bacteria (Wan et al., 2024). We highlight in Table 1 the comparison between the second peptide with the other three peptides. B1-B7 are Gram negative bacteria, while B8-B11 are Gram positive. Peptide 2 specialized to the Gram positive bacteria (B8-B11 scores predicted highly active compared to the other peptides) at the expense of broad spectrum activity for Gram negatives (B5, B6, B7 predictied inactive). By specializing its solutions to Gram negative and Gram positive bacteria separately, MOCOBO achieves low MIC across the target bacteria with only $K=4$ peptides. A similar table of results is also provided for the template constrained peptide design task (see Table A.1).
In Figure 4, we provide in vitro results for the two best “template free” and two best “template constrained” runs of MOCOBO for the peptide design task. Here, “best” means runs that achieved highest coverage scores according to the APEX 1.1 model. For each run, Figure 4 provides the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target pathogenic bacteria. Results demonstrate that solutions found by MOCOBO optimizing against the in-silico APEX 1.1 model achieve good coverage of the $11$ target pathogenic bacteria in vitro. In Figure A.1, we provide the full set of in vitro results for each these runs of MOCOBO, with MIC values obtained by each of the $K=4$ peptides found for each target bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
@@
In the four leftmost panels of Figure 5, we depict the $T=4$ obstacle courses that the rover seeks to navigate with a covering set of $K=2$ optimized trajectories (magenta, blue). Together, the two optimized trajectories cover the $T=4$ obstacle courses such that all obstacles are avoided. In Figure A.3 and A.4, we provide analogous figures for the $T=8$ and $T=12$ variations of the rover task. In each example, the MOCOBO optimized set of trajectories covers all obstacle courses such that all obstacles are avoided.
@@
In the Figure 5 (five rightmost panels) and Figure A.2, we provide the original hdr images and the images obtained by the $K=4$ solutions found by a single run of MOCOBO for the church and desk test images respectively. In both variations, three of the four solutions result in high-quality images, while one solution results in a poor quality image (see middle right church image in Figure A.2 and rightmost desk image in Figure 5). The two poor quality images were both generated by the one solution in the best covering set that MOCOBO used to cover metrics $5$ and $6$ (see metric ID numbers in Table A.3). This indicates that metrics $5$ and $6$ are poor indicators of true image quality in this setting. Overall, this result highlights a useful application of MOCOBO: By dedicating one of the $K$ solutions to maximizing the misleading metrics ($5$ and $6$), the other $3$ solutions to can focus on the remaining subjectively better quality metrics.
A notable apparent limitation of the existing IAA/IQA metrics we used here is that they all favored monochromatic images for both the church and desk images. As such, the tone-mapped images obtained appear less colorful than their hand-tuned counterparts reported in prior work.

### v2

@@ 3.1.2 Greedy approximation of $S^{*}_{D_{s}}$
0:  Dataset $D_{s}$, observed values $f_{t}(\mathbf{x})$ for all $t\in\{1,\ldots,T\}$ and $\mathbf{x}\in D_{s}$, set size $K$.
  Initialize $A\leftarrow\emptyset$. {Start with an empty set.}
  Compute the initial coverage score for $A$: $c(A)\leftarrow 0$.
  for $k=1$ to $K$ do
     Initialize $\mathbf{x}_{\text{best}}\leftarrow\text{None}$ and $\Delta_{\text{best}}\leftarrow-\infty$.
     for $\mathbf{x}\in D_{s}\setminus A$ do
        Compute marginal coverage of adding $\mathbf{x}$ to $A$:
@@
        if $\Delta c>\Delta_{\text{best}}$ then
           $\mathbf{x}_{\text{best}}\leftarrow\mathbf{x}$ and $\Delta_{\text{best}}\leftarrow\Delta c$.
        end if
     end for
     Update $A\leftarrow A\cup\{\mathbf{x}_{\text{best}}\}$, $c(A)\leftarrow c(A)+\Delta_{\text{best}}$.
  end for
  Output: $A^{*}_{D_{s}}\leftarrow A$.
Since constructing $S^{*}_{D_{s}}$ is NP-hard, we use an approximate construction of $S^{*}_{D_{s}}$ on each step of optimization $s$. We present Algorithm 1 with $O(nKT)$ running time that provides a $(1-\frac{1}{e})$-approximation of $S^{*}_{D_{s}}$.
Theorem 3.2.
The set $A^{*}_{D_{s}}$ output by Algorithm 1 satisfies $c(A^{*}_{D_{s}})\geq\left(1-\frac{1}{e}\right)c(S^{*}_{D_{s}})$.
@@
The algorithm iterates $K$ times to construct the covering set. In each iteration, it evaluates at most $n$ candidate points, and for each candidate, it computes the incremental coverage score by iterating over $T$ objectives. The runtime is thus $O(K\cdot n\cdot T)$. For practical applications where we can assume relatively small $K$ and $T$, the runtime is approximately $O(n)$. For an empirical evaluation of the wall-clock runtime, see Section A.2.3.
@@
This follows from the fact that we reduced from Max $k$-cover, for which a better approximation ratio is not practically achievable unless $P=NP$ [19]. ∎
@@ 3.1.3 Extending ECI to the batch acquisition setting (q-ECI)
In batch acquisition, we select a batch of $q>1$ candidates for evaluation. Following recent work on batch EI [71, 67], we define q-ECI, a natural extension of ECI to the batch setting:
$q\text{-ECI}(\mathbf{X})$ $=\mathbb{E}_{p(\mathbf{Y}|\mathbf{X},D)}\left[\max(0,\max_{r=1,\dots,q}c\left(S^{*}_{D_{s}\cup\left\{(\mathbf{x}_{r},\mathbf{y}_{r})\right\}}\right)-c\left(S^{*}_{D_{s}})\right)\right]$ (6)
q-ECI gives the expected improvement in coverage score after simultaneously observing the batch of $q$ points $\mathbf{X}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{q}\right\}$. However, the resulting Monte Carlo expectation would require $O(q\times m)$ evaluations of Algorithm 1. We therefore adopt a more approximate batching strategy for practical use with large $q$ (see Section A.3 for details).
3.1.4 MOCOBO trust regions
In order to find a set of $K$ solutions, MOCOBO maintains $K$ simultaneous local optimization runs using $K$ individual trust regions. Each local run $k\in\left\{1,\ldots,K\right\}$ aims to find a single solution $\mathbf{x}^{*}_{k}$, which together form the desired set $S^{*}$. As in the original TuRBO paper, trust regions are rectangular regions of the search space $\mathcal{T}_{k}\subseteq\mathcal{X}$ defined solely by their size and center point. We center the $K$ trust regions on the best covering set of solutions observed so far during optimization. In particular, on each optimization step $s$, we construct an approximation of $S^{*}_{D_{s}}$ using Algorithm 1, and center each trust region $\mathcal{T}_{k}$ on the corresponding point $\mathbf{x}^{*(s)}_{k}$ in $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\right\}$. We then use our proposed ECI (or q-ECI when $q>1$) acquisition function to select $q$ candidates for evaluation from within each trust region. The MC approximation of ECI described in Section 3.1.1 can be straightforwardly applied to select candidates in trust region $\mathcal{T}_{k}$ by sampling the $m$ discrete points from within the rectangular bounds of $\mathcal{T}_{k}$. Since we select and evaluate $q$ candidates from each of the $K$ trust regions, the total number of observed data points $n$ at step $s$ of MOCOBO is $n=s\times K\times q$. As in the original TuRBO algorithm, each trust region $\mathcal{T}_{k}$ has success and failure counters that dictate the size of the trust region. For MOCOBO, we count a success for trust region $\mathcal{T}_{k}$ whenever $\mathcal{T}_{k}$ proposes a candidate on step $s$ that improves upon the best coverage score and is included in $S^{*}_{D_{s+1}}$.
@@ 4 Experiments
We implement MOCOBO using BoTorch [2] and GPyTorch [20]. Code to reproduce all results in the paper is available on GitHub: placeholder. We use an acquisition batch size of $20$ for all tasks and across all BO methods compared. Further implementation details are provided in subsection A.7. For our structured optimization tasks, we apply MOCOBO in the latent space of a pre-trained VAE on which we perform regular end-to-end updates with the surrogate model during optimization [44].
@@
In Figure 2, we plot the best coverage score $c$ Equation 1 obtained by the $K$ best covering solutions found so far after a certain number of function evaluations on each task. Since BO baseline methods are not designed to optimize coverage directly, instead aiming to find a single solution or a set of solutions to optimize the $T$ objectives, we plot the best coverage score $c$ obtained by the $K$ best covering solutions found by the method. All plots show mean coverage scores averaged over $20$ replications of each method, and show standard errors.
Baselines.
In all plots, we compare MOCOBO against CluSO, TuRBO, ROBOT, and MORBO [42, 17, 45, 11]. We also compare to sampling uniformly at random in the search space. For each method, we compare to the performance of the best set of $K$ solutions found.
@@
In the case of the two structured optimization tasks (molecule and peptide design), we apply LOL-BO and ROBOT as described by design for these problem settings. For MORBO and CluSO, we take the straightforward approach of applying the methods directly in the continuous latent space of the VAE model without other adaptations. The same pre-trained VAE model is used across methods compared.
Table 1: Best set of $K=4$ peptides found by one run of MOCOBO for the “template free" peptide design task. We provide the APEX model’s predicted MIC for each sequence on each of the $11$ target pathogenic bacteria. The target pathogenic bacteria B1,,$\ldots$,, … , B11 are listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive pathogenic bacteria respectively. The best/lowest MIC achieved for each pathogenic bacteria is in bold in each column. See row “TF1" in Figure 3 left, and Figure A.1 for in vitro MICs for this set of $K=4$ peptides. See Table A.1 for an analogous result on the “template constrained" peptide design task.
Extending baselines to the coverage optimization setting.
As TuRBO, LOL-BO, and ROBOT target single-objective optimization, we conduct $T$ independent runs to optimize each of the $T$ objectives for the multi-objective task. We use all solutions gathered from the $T$ runs to compute the best covering set of $K$ solutions found by each method. Unlike LOL-BO and TuRBO which seek a single best solution for a given objective, a single run of ROBOT seeks a set of $M$ solutions that are pairwise diverse. We run ROBOT with $M=K$ so that each independent run of ROBOT seeks $M=K$ diverse solutions. The aggregate result of the $T$ independent runs for ROBOT is thus $K$ diverse solutions for each of the $T$ objectives. We compare to the best covering $K$ solutions from among those $K*T$ solutions. See Section A.7.4 for more details on the diversity constraints used by ROBOT and the associated hyperparameters. MORBO is a multi-objective optimization method and can thus be applied directly to each multi-objective optimization task. For each run of MORBO, we compare to the best covering set of $K$ solutions found among all solutions proposed by the run. For CluSO, no extension is needed as this method is designed for coverage optimization. We run CluSO with the same $K$ used by MOCOBO, and all other hyperparameters set to CluSO defaults.
@@ 4.1 Tasks
We evaluate MOCOBO on two variations of the peptide design task: “template free" (TF) and “template constrained" (TC). For TF, we allow the optimizer to propose any sequence of amino acids. For TC, we add a constraint that any sequence proposed by the optimizer must have a minimum of $75$ percent sequence similarity to at least one of the $10$ template amino acid sequences in Table A.8. These $10$ templates were mined from extinct organisms and selected by Wan et al. [66]. The motivation of the template constrained task is to design peptides specifically likely to evade antibiotic resistance by producing “extinct-like” peptides that bacteria have not encountered in nature in thousands of years. For MOCOBO and all BO baselines (TuRBO, LOL-BO, ROBOT, and MORBO), we handle the optimization constraint by adapting techniques from SCBO [16]. For CluSO and random sampling, we handle the constraint with rejection sampling.
@@
Ranolazine is a drug used to treat chest pain. The original Ranolazine MPO task from the Guacamol benchmark suite of molecular design tasks [6] aims to design an alternative to this drug: a molecule with a high fingerprint similarity to Ranolaize that includes fluorine. We extend this task to the multi-objective optimization setting by searching for alternatives to Ranolazine that include $T=6$ reactive nonmetal elements not found in Ranolazine: fluorine, chlorine, bromine, selenium, sulfur, and phosphorus. We aim to cover the $T=6$ objectives with $K=3$ molecules. We use the SELFIES-VAE introduced by Maus et al. [44] to map the molecular space to a continuous $256$ dimensional space.
Figure 3: (Left) In vitro results for the two best “template free" (TF1, TF2) and two best “template constrained" (TC1, TC2) runs of MOCOBO for the peptide design task. Columns are the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target pathogenic bacteria B1,,$\ldots$,, … , B11 listed in Table A.7. (-) and (+) indicate Gram negative and Gram positive respectively. TF1 and TC1 correspond to the single runs of MOCOBO shown in Table 1 and Table A.1 respectively. Methods used to obtain in vitro MICs are provided in Section A.1. (Right) Ablation study comparing MOCOBO to optimization performance where a known “good” partitioning of the $T$ objectives into $K$ subsets is available in advance. We individually optimize $K$ solutions, one for each partition.
@@
The rover optimization task introduced by Wang et al. [68] consists of finding a $60$-dimensional policy that allows a rover to move along some trajectory while avoiding a fixed set of obstacles. To frame this as a multi-objective optimization task, we design $T$ unique obstacle courses for the rover to navigate. The obstacle courses are designed such that no single policy can successfully navigate all courses. We seek $K<T$ policies so that at least one policy enables the rover to avoid obstacles in each course. We evaluate on three unique instances with varying numbers of obstacle courses. For the instances of this task with $T=4$ and $T=8$, we seek to cover the objectives with $K=2$ solutions. For the $T=12$ instance, we seek to cover the objectives with $K=4$ solutions.
@@
In high dynamic range (HDR) images, some pixels (often associated with light sources) can dominate overall contrast, requiring adjustments to reveal detail in low-contrast areas, a problem known as tone mapping [51, Section 6]. Tone mapping algorithms involve various tunable parameters, resulting in a high-dimensional optimization problem of subjectively perceived quality. We seek a covering set of $K=4$ solutions to optimize a set of $T=7$ image aesthetic (IAA) and quality (IQA) assessment metrics from the pyiqa library [9] (see metrics listed in Table A.5). Our practical goal is that, while we do not know a priori which metric is best for a particular image, covering all metrics may result in at least one high quality image. We optimize over the $13$-dimensional parameter space of an established tone mapping pipeline to tone-map the “Stanford Memorial Church" [12] and “desk lamp" [7] benchmark images. See Section A.8.1 for details.
@@ 4.2 Optimization results
In Figure 2, we provide optimization results comparing MOCOBO to the baselines discussed above on all tasks. The results show that MOCOBO finds sets of $K$ solutions that achieve higher coverage scores across tasks. The “T Individually Optimized Solutions" baseline appears as a horizontal dotted line in all plots of Figure 2, representing the average coverage score of $T$ individually optimized solutions, serving as an approximation of the best possible performance without the constraint of a limited $K<T$ solution set. Results in Figure 2 demonstrate that the smaller set of $K<T$ solutions identified by MOCOBO nearly equals the performance of the complete set of $T$ individually optimized solutions. This result depends on using domain knowledge to choose $K$ large enough to achieve it. Results in Figure 2 show that matching the performance of $T$ optimized solutions is possible with some values of $K\ll T$.
@@
In Table 1, we provide the $K=4$ peptides found by one run of MOCOBO for the template free (TF) peptide design task. For each peptide, we provide the the APEX 1.1 model’s predicted MIC for each of the $11$ target bacteria. MIC values $\leq 16$ $\mu$ mol $L^{-1}$ are considered to be “highly active" against the target pathogenic bacteria [66]. We highlight in Table 1 the comparison between the second peptide with the other three peptides. B1-B7 are Gram negative (GN) bacteria, while B8-B11 are Gram positive (GP). Peptide 2 specialized to the GP bacteria (B8-B11 scores predicted highly active compared to the other peptides) at the expense of broad spectrum activity for the GN bacteria (B5, B6, B7 predicted inactive). By specializing its solutions to GN and GP bacteria separately, MOCOBO achieves low MIC across the target bacteria with only $K=4$ peptides. A similar table of results is also provided for the template constrained (TC) peptide design task (see Table A.1).
Figure 4: (Four Leftmost Panels) $T=4$ varied obstacle courses as multiple objectives for the rover task and $K=2$ covering trajectories (blue, magenta) found by MOCOBO. The first trajectory (magenta) successfully navigates obstacle courses 2 and 3 (second and third panels from the left). The second trajectory (blue) navigates obstacle courses 1 and 4 (first and fourth panels from the left). (Middle Panel) Initial hdr desk image for the desk variation of the “image tone mapping" task. (Four Rightmost Panels) Images obtained by transforming the hdr desk image using the best covering set of $K=4$ solutions found by a single run of MOCOBO.
In Figure 3 left, we provide in vitro results for the two best TF and two best TC runs of MOCOBO for the peptide design task. Here, “best" means runs that achieved highest coverage scores according to the APEX 1.1 model. For each run, Figure 3 left provides the best/lowest in vitro MIC among the $K=4$ peptides found by MOCOBO for each target bacteria. Results demonstrate that solutions found by MOCOBO optimizing against the in-silico APEX 1.1 model achieve good coverage of the $11$ target pathogenic bacteria in vitro. In Figure A.1, we provide the full set of in vitro results for each these runs of MOCOBO, with MIC values obtained by each of the $K=4$ peptides found for each target bacteria. Methods used to obtain in vitro MICs are provided in Section A.1.
@@
In Figure 4 left, we depict the $4$ varied obstacle courses used for the $T=4$ variation of the rover task and $K=2$ covering trajectories found by a single run of MOCOBO. In Figure A.3 and A.4, we provide analogous figures for the $T=8$ and $T=12$ variations. In each example, the MOCOBO optimized set of trajectories covers all obstacle courses such that all obstacles are avoided.
@@
In Figure 4 right and Figure A.2, we provide the original hdr images and the images obtained by the $K=4$ solutions found by a single run of MOCOBO for the church and desk test images respectively. In both variations, three of the four solutions result in high quality images, while one results in a poor quality image (see middle right church image and rightmost desk image). The poor quality images were each generated by the one solution that MOCOBO used to cover metrics $5$ and $6$ (see metric ID numbers in Table A.5). This indicates that metrics $5$ and $6$ are poor indicators of true image quality in this setting. This result highlights a useful application of MOCOBO: By dedicating one of the $K$ solutions to maximizing the misleading metrics ($5$ and $6$), the other three solutions to can focus on the remaining subjectively better metrics.